# Continuum Limits of Lazy Open Quantum Walks

### v1

@@ Abstract
We derive the continuous spacetime limit of the one dimensional lazy discrete time quantum walk, obtaining explicit macroscopic evolution equations for a three state model in the presence of decoherence. While continuum limits of two state quantum walks are well established, an explicit continuous spacetime formulation for the lazy three state walk, particularly including noise, has not previously been constructed. Using an $\mathcal{SU}$(3) representation of a Grover type coin together with a Lindblad formulation of decoherence acting either on the coin or the spatial subspace, we systematically expand the discrete dynamics in both space and time to obtain continuum master equations governing the coarse grained evolution. The resulting generators yield a genuine partial differential equation description of the walk, going beyond purely probabilistic or spectral correspondences. We show that the unitary limit is governed by a Dirac-type SU(3) Hamiltonian describing ballistic advection of left and right moving modes coupled by local symmetric mixing, with the rest state acting as an additional internal degree of freedom. Coin dephasing selectively damps internal coherences while preserving coherent spatial transport, whereas spatial dephasing suppresses long range spatial interference and rapidly drives the dynamics toward classical behaviour. This continuum framework clarifies how internal symmetry, rest state coupling, and distinct decoherence channels shape large scale transport in lazy open quantum walks, and provides a foundation for future extensions toward multichannel quantum transport models and quantum-inspired algorithms.
Journal Title
@@ 1 Introduction
Quantum walks constitute the quantum analogue of classical random walks and form a versatile framework for modelling quantum transport and computation [Ambainis, 2004, Molfetta, 2021]. Unlike their classical counterparts, quantum walks exhibit interference effects that can lead to quadratic speedups in hitting and mixing times, underpinning their use in quantum algorithms for search and graph analysis [Kempe, 2002, 2003, Ambainis, 2007]. Both continuous time and discrete time formulations have been developed [Farhi and Gutmann, 1998, Aharonov et al., 1993], with discrete time quantum walks (DTQWs) serving as the most direct analogue of lattice based classical walks.
@@
In the discrete time setting, the walker occupies a position register with an internal coin space, whose rotation governs directional propagation. The variant relevant for this work is the lazy quantum walk, originally introduced by Childs [2009], to demonstrate that the continuous time quantum walk arises as a limit of the discrete model. The inclusion of a rest state introduces a non-zero probability of remaining at the current site, modifying the spectrum of the walk operator and enabling a smooth connection between discrete and continuous time dynamics. Unlike the approach of Childs [2009], where the lazy walk was introduced to embed continuous time dynamics at the spectral level, our analysis presented here includes dephasing, and derives an explicit continuous spacetime master equation by expanding the discrete Lindblad map in space and time. This yields a genuine coarse grained partial differential equation (PDE) description rather than a spectral correspondence between discrete and continuous models.
@@
Beyond foundational interest, lazy walks have also been employed in algorithmic contexts, such as Grover’s search, where self loops at vertices act as effective rest states, influencing success probabilities [Wong, 2015, 2018, Giri and Korepin, 2019]. Including a rest state extends the walk to a three channel model that more closely resembles discrete velocity schemes in kinetic theory [Childs, 2009, Succi, 2018]. The stationary component provides a minimal setting where propagating and stationary dynamics coexist [Molfetta and Debbasch, 2012]. This may therefore serve as a foundation for future developments using open quantum walks to model coarse grained transport or hydrodynamic behaviour. Continuous spacetime limits provide a bridge between discrete quantum walks and continuum physical theories [Manighalam and Kon, 2020, Molfetta and Arrighi, 2019]. By expanding the discrete step operator in powers of a small parameter that controls the time and lattice spacing, one can identify the effective PDE governing large scale dynamics, ranging from Dirac-type to dispersive behaviour, depending on the chosen coin and scaling [Jolly and Di Molfetta, 2023, Yamagishi et al., 2023]. This approach makes it possible to interpret microscopic unitary rules as discrete simulators of macroscopic behaviour.
@@
Open quantum walks (OQWs) generalise the framework to include environmental noise or engineered dissipation through completely positive trace preserving (CPTP) maps. Environment induced errors in discrete quantum walks were first studied by Dür et al. [2002] to model errors in proposed experimental implementations. Kendon and Tregenna [2003] found that decoherence could improve the algorithmic properties of quantum walks by providing a smoother uniform sampling that was still quadratically faster than classical. Dissipation has been leveraged more generally for quantum computing and quantum state preparation [Verstraete et al., 2009]. More general open system quantum walk models were introduced by Attal et al. [2012a, b], Sinayskiy and Petruccione [2012]. They provide a natural setting for describing transport in open quantum systems and have found applications in dissipative computing, state transfer, and biological energy transport [Mohseni et al., 2008, Sinayskiy and Petruccione, 2014].
@@
In this paper, we analyse a lazy DTQW subject to two distinct noise channels: one acting on the coin degree of freedom and another on the spatial register. By performing a systematic scaling analysis of the discrete evolution, we derive the corresponding continuum master equation governing the coarse grained state. This provides insight into the behaviour of lazy quantum walks with decoherence and complements existing probabilistic approaches, which establish Gaussian limits but not explicit macroscopic dynamics. For completeness, we derive the continuum limit for both the coin and spatial decoherence models. However, previous studies have shown that coin dephasing preserves spatial coherence while suppressing only internal superpositions, making it more useful for modelling controlled decoherence in quantum transport and algorithmic settings [López and Paz, 2003, Love and Boghosian, 2005, Kendon and Tregenna, 2003]. Here we present the continuous spacetime limit of the lazy three state walk in a full $\mathrm{SU}(3)$ representation, where the unitary and dephasing generators emerge naturally from the Gell Mann structure. This offers a clear route to continuum models for multichannel quantum walks and highlights how internal symmetry and coupling shape large scale dynamics.
@@ 2 Lazy Discrete Time Quantum Walks
where $|u\rangle=\tfrac{1}{\sqrt{3}}(1,1,1)^{\mathsf{T}}$, which is as unbiased as possible across all three outcomes [Saha et al., 2021]. The Grover coin provides a simple and highly symmetric choice for the internal dynamics of the walk. Different coin parametrisations would, in general, lead to distinct effective behaviours in the continuum limit. In the present work, we focus on the symmetric Grover-type coin in order to preserve translational invariance and maintain a balanced coupling between all internal states. This choice ensures that any emergent large scale dynamics arise solely from the interplay between coherent evolution and decoherence, rather than from asymmetries in the underlying coin operation. Nonetheless, the symmetry can be intentionally broken if one wishes to introduce directional bias or model anisotropic effects.
@@ 2.1 Fourier Basis
In order to analyse the continuum limit, it is convenient to work in the momentum representation, where the shift operator becomes diagonal and separable from the coin operator [Ambainis et al., 2001]. We therefore first transform both operators to Fourier space, before expressing them in the $\mathrm{SU}(3)$ Gell–Mann basis used for the continuum expansion. To obtain the momentum–space form of the walk, we apply the spatial Fourier transform:
@@
$\hat{S}(k)=e^{-ik\Delta_{x}}|L\rangle\langle L|\;+\;|S\rangle\langle S|\;+\;e^{+ik\Delta_{x}}|R\rangle\langle R|$ (7)
@@
$\tilde{\boldsymbol{\psi}}(k,t{+}1)=\mathbf{\hat{S}}(k)\,\mathbf{\hat{C}}_{G}\,\tilde{\boldsymbol{\psi}}(k,t).$ (8)
@@
$\hat{S}(k)=\begin{pmatrix}e^{-ik\Delta_{x}}&0&0\\[4.0pt] 0&1&0\\[4.0pt] 0&0&e^{ik\Delta_{x}}\end{pmatrix}$ (9)
which shows explicitly that the $|R\rangle$ component shifts forward, the $|L\rangle$ component shifts backward, and the $|S\rangle$ component remains localised. The Grover coin, given by equation (3) admits a simple spectral decomposition, which enables us to parameterise it as the exponential of a generator. We define $G=\mathbb{I}_{3}-|u\rangle\langle u|$ where $|u\rangle=\tfrac{1}{\sqrt{3}}(1,1,1)^{\mathsf{T}}$, then:
@@
A convenient parametrisation is given by (See Appendix A for derivation of G):
$\mathbf{\hat{C}}(\theta)=e^{-i\theta G}=e^{-i\theta(\mathbb{I}_{3}-|u\rangle\langle u|)}=|u\rangle\langle u|+e^{-i\theta}\,(\mathbb{I}_{3}-|u\rangle\langle u|)$ (11)
In this representation, the eigenstate corresponding to the uniform vector $|u\rangle=\tfrac{1}{\sqrt{3}}(1,1,1)^{\mathsf{T}}$ has eigenvalue +1+1, while the two orthogonal Fourier modes acquire a common phase $e^{-i\theta}$. Thus, the family interpolates smoothly around the Grover operator, which corresponds to the special value $\theta=\pi$. This parametrisation introduces a tunable coin angle that controls the relative phase between the symmetric mode and the two transverse modes, and is therefore the natural small parameter for the continuum expansion. The parameter $\theta$ controls the strength of mixing between the three coin states. In the continuum limit this coupling plays the role of an effective mass term. Small $\theta$ yields nearly ballistic propagation, while larger values increase local mixing and open a mass-like gap in the dispersion. The standard Grover coin then corresponds to the special rotation angle $\theta=\pi$. The full walk operator in momentum space can be written as:
$\mathbf{\hat{U}}(k,\theta)=\mathbf{\hat{S}}(k)\,\mathbf{\mathbf{\hat{C}}}(\theta)$ (12)
@@ 2.2 Gell Mann Representation
We now express $\hat{S}(k)$ and $\mathbf{\hat{C}}(\theta)$ in the $\mathcal{SU}$(3) Gell Mann basis. This representation makes the underlying generators explicit and is the natural starting point for performing the continuum limit expansion of the walk operator in section 3.
For two state quantum walks, the two-dimensional coin operators can be decomposed in the Pauli basis, any $\mathcal{SU}$(2) operator can be written as a linear combination of Pauli matrices. In the lazy quantum walk, the coin space is three dimensional, so the natural generalisation is to use the eight Gell Mann matrices $\{\lambda_{i}\}_{i=1}^{8}$, traceless Hermitian operators, which, together with the identity $\mathbb{I}_{3}$, span the space of $3\times 3$ Hermitian operators [Wilson, 2024]. The explicit form of each Gell Mann matrix can be found in Appendix B. They satisfy the orthogonality relation $\mathrm{Tr}(\lambda_{i}\lambda_{j})=2\delta_{ij}$, so that any $\mathcal{SU}(3)$ operator $A\in\mathbb{C}^{3\times 3}$ can be expanded as
@@
Having expressed both the shift and coin in diagonal form in Fourier space, we now rewrite these operators in the $\mathrm{SU}(3)$ Gell Mann basis. We first define the $\mathcal{SU}(3)$ generator:
$J_{z}^{(3)}\;=\;\tfrac{1}{2}\lambda_{3}+\tfrac{\sqrt{3}}{2}\lambda_{8}\;=\;\mathrm{diag}(1,0,-1)$ (14)
@@
$\mathbf{\hat{S}}(k)=\exp\!\bigl[-\,i\,\Delta_{x}k\,(\mathbb{I}_{x}\otimes J_{z}^{(3)})\bigr]$ (15)
@@
which mixes the three internal states with equal weight.
@@ 2.3 Decoherence in Quantum Walks
Decoherence is usually described as a non-unitary evolution using the density matrix formalism to describe the mixed states that arise. Simple decoherence models for quantum walks were introduced in [Kendon and Tregenna, 2003]. Here, we use a similar simple Lindblad master equation to examine the basic features of decoherence in the lazy quantum walk, providing a baseline against which more complex decoherence mechanisms can be compared. This approach assumes a Markovian open system description, in which the environment has a short correlation time and no memory effects are retained. The Lindblad master equations provides dynamics that are completely positive and trace preserving, ensuring that the density operator remains physical at all times, regardless of the step size as we take the limits [Annabestani et al., 2010]. The coupling to the environment is treated phenomenologically through a single decoherence strength $\gamma$, without specifying a microscopic bath model or spectral density. Initial system-environment correlations are neglected, and the system follows a weak decoherence regime [Breuer and Petruccione, 2002].
All the operators act in the same Hilbert space $\mathcal{H}_{xc}=\mathcal{H}_{x}\otimes\mathcal{H}_{c}$ as the unitary quantum walk. The system is described by a time dependent density operator, expressed in the computational basis, as:
$\boldsymbol{\hat{\rho}}(t)=\sum_{x,c}\sum_{x^{\prime},c^{\prime}}\boldsymbol{\hat{\rho}}_{x,c}^{x^{\prime},c^{\prime}}(t)|x,c\rangle\langle x^{\prime},c^{\prime}|.$ (17)
@@
$\boldsymbol{\hat{\rho}}(t+\Delta_{t})=(1-\gamma)\hat{\mathbf{U}}\boldsymbol{\hat{\rho}}(t)\hat{\mathbf{U}}^{\dagger}+\gamma\sum_{j}\mathbb{P}_{j}\hat{\mathbf{U}}\boldsymbol{\hat{\rho}}(t)\hat{\mathbf{U}}^{\dagger}\mathbb{P}_{j}^{\dagger}$ (18)
where $\gamma\in[0,1]$ is the probability of decoherence per full step, and the operator $\hat{\mathbf{U}}$ is the full unitary evolution operator. This decoherence model was reviewed in Kendon [2007] and evolved numerically for various choices of $\mathbb{P}_{j}$. This is the infinite temperature Lindblad form, whose stationary state is the maximally mixed state and treats all energy levels symmetrically.
@@
The decoherence is modelled as projective noise using jump operators are constructed as
@@
Figure 2 compares the effects of coin and spatial dephasing on the lazy DTQW for several values of the decoherence strength $\gamma$ and for different coin angles $\theta$. Coin dephasing, shown in panels (a,c), selectively damps the off-diagonal coherences between the internal coin states $\{|L\rangle,|S\rangle,|R\rangle\}$ while leaving spatial coherences intact. As a result, the ballistic envelope characteristic of the unitary lazy walk is preserved, although the interference fringes decorating the left, rest, and right moving channels are progressively suppressed as $\gamma$ increases. The three peak structure therefore remains visible over a broad range of $\gamma$. In contrast, spatial dephasing, shown in panels (b,d), suppresses off-diagonal density matrix elements $\rho(x,x^{\prime})$ for $x\neq x^{\prime}$, thereby directly removing the long range spatial interference responsible for quantum ballistic transport. As $\gamma$ increases, the ballistic side peaks collapse more rapidly than in the coin dephased case and the probability distribution evolves toward a broad, approximately Gaussian profile centred at the origin. This transition occurs already at relatively small values of $\gamma$, reflecting the sensitivity of ballistic transport to spatial coherence. In the limit $\gamma\to 1$, both decoherence mechanisms yield a similar Gaussian distribution, corresponding to a fully classical diffusive regime. Comparing different coin angles further reveals the role of the coin in controlling transport and localisation. For $\theta=\pi$, the unitary walk exhibits a pronounced central peak, indicating that the coin strongly favours population of the rest state and suppresses propagation. For $\theta=\pi/2$, the coin induces stronger mixing between the propagating and non-propagating components, resulting in a less pronounced central peak and a higher probability of reaching the ballistic fronts. This behaviour persists under both decoherence mechanisms, although spatial dephasing again accelerates the transition toward a Gaussian profile. Overall, increasing $\theta$ modifies the internal coin mixing and interference structure of the walk, leading to systematically reduced spatial spreading and enhanced localisation near the origin.

### v2

@@ Abstract
We derive the continuous spacetime limit of the one dimensional lazy discrete time quantum walk, obtaining explicit macroscopic evolution equations for a three state model in the presence of decoherence. While continuum limits of two state quantum walks are well established, an explicit continuous spacetime formulation for the lazy three state walk, particularly including noise, has not previously been constructed. Using an $\mathcal{SU}$(3) representation of a Grover type coin together with a Lindblad formulation of decoherence acting either on the coin or the spatial subspace, we systematically expand the discrete dynamics in both space and time to obtain continuum master equations governing the coarse grained evolution. The resulting generators yield a genuine partial differential equation description of the walk, going beyond purely probabilistic or spectral correspondences. We show that the unitary limit is governed by a Dirac-type $\mathcal{SU}(3)$ Hamiltonian describing ballistic advection of left and right moving modes coupled by local symmetric mixing, with the rest state acting as an additional internal degree of freedom. Coin dephasing selectively damps internal coherences while preserving coherent spatial transport, whereas spatial dephasing suppresses long range spatial interference and rapidly drives the dynamics toward classical behaviour. This continuum framework clarifies how internal symmetry, rest state coupling, and distinct decoherence channels shape large scale transport in lazy open quantum walks, and provides a foundation for future extensions toward multichannel quantum transport models and quantum inspired algorithms.
Paper
@@ 1 Introduction
Quantum walks constitute the quantum analogue of classical random walks and form a versatile framework for modelling quantum transport and computation [Kendon, 2020, Molfetta, 2021]. Unlike their classical counterparts, quantum walks exhibit interference effects that can lead to quadratic speedups in hitting and mixing times, underpinning their use in quantum algorithms for search and graph analysis [Shenvi et al., 2003, Ambainis, 2004, 2007]. Both continuous time and discrete time formulations have been developed [Farhi and Gutmann, 1998, Aharonov et al., 1993], with discrete time quantum walks (DTQWs) serving as the most direct analogue of lattice based classical walks.
@@
In the discrete time setting, the walker occupies a position register with an internal coin space, whose rotation governs directional propagation. The variant relevant for this work is the lazy quantum walk, originally introduced by Childs [2009] to demonstrate that the continuous time quantum walk arises as a limit of the discrete model. The inclusion of a rest state introduces a non-zero probability of remaining at the current site, modifying the spectrum of the walk operator and enabling a smooth connection between discrete and continuous time dynamics. Unlike the approach of Childs [2009], where the lazy walk was introduced to embed continuous time dynamics at the spectral level, our analysis presented here includes dephasing, and derives an explicit continuous spacetime master equation by expanding the discrete Lindblad map in space and time. This yields a genuine coarse grained partial differential equation (PDE) description rather than a spectral correspondence between discrete and continuous models.
@@
Beyond foundational interest, lazy walks have also been employed in algorithmic contexts, such as Grover’s search, where self loops at vertices act as effective rest states, influencing success probabilities [Wong, 2015, 2018, Giri and Korepin, 2019]. Including a rest state extends the walk to a three channel model that more closely resembles discrete velocity schemes in kinetic theory [Succi, 2018]. The stationary component provides a minimal setting where propagating and stationary dynamics coexist [Molfetta and Debbasch, 2012]. This may therefore serve as a foundation for future developments using open quantum walks to model coarse grained transport or hydrodynamic behaviour. Continuous spacetime limits provide a bridge between discrete quantum walks and continuum physical theories [Manighalam and Kon, 2020, Molfetta and Arrighi, 2019]. By expanding the discrete step operator in powers of a small parameter that controls the time and lattice spacing, one can identify the effective PDE governing large scale dynamics, ranging from Dirac-type to dispersive behaviour, depending on the chosen coin and scaling [Jolly and Di Molfetta, 2023, Yamagishi et al., 2023]. This approach makes it possible to interpret microscopic unitary rules as discrete simulators of macroscopic behaviour.
@@
Open quantum walks (OQWs) generalise the framework to include environmental noise or engineered dissipation through completely positive trace preserving (CPTP) maps. Environment induced errors in discrete quantum walks were first studied by Dür et al. [2002] to model errors in proposed experimental implementations. Kendon and Tregenna [2003] found that decoherence could improve the algorithmic properties of quantum walks by providing a smoother uniform sampling that was still quadratically faster than classical. Dissipation has been leveraged more generally for quantum computing and quantum state preparation [Beige et al., 2000, Verstraete et al., 2009]. More general open system quantum walk models were introduced by Attal et al. [2012a, b], Sinayskiy and Petruccione [2012]. They provide a natural setting for describing transport in open quantum systems and have found applications in dissipative computing, state transfer, and biological energy transport [Mohseni et al., 2008, Sinayskiy and Petruccione, 2014].
@@
In this paper, we analyse a lazy DTQW subject to two distinct noise channels: one acting on the coin degree of freedom and another on the spatial register. By performing a systematic scaling analysis of the discrete evolution, we derive the corresponding continuum master equation governing the coarse grained state. This provides insight into the behaviour of lazy quantum walks with decoherence and complements existing probabilistic approaches, which establish Gaussian limits but not explicit macroscopic dynamics. For completeness, we derive the continuum limit for both the coin and spatial decoherence models. However, previous studies have shown that coin dephasing preserves spatial coherence while suppressing only internal superpositions, making it more useful for modelling controlled decoherence in quantum transport and algorithmic settings [López and Paz, 2003, Love and Boghosian, 2005, Kendon and Tregenna, 2003]. Here we present the continuous spacetime limit of the lazy three state walk in a full $\mathcal{SU}(3)$ representation, where the unitary and dephasing generators emerge naturally from the Gell Mann structure. This offers a clear route to continuum models for multichannel quantum walks and highlights how internal symmetry and coupling shape large scale dynamics.
@@ 2 Lazy Discrete Time Quantum Walks
where
$|u\rangle=\tfrac{1}{\sqrt{3}}(1,1,1)^{\mathsf{T}}.$ (4)
The coin operator is as unbiased as possible across all three outcomes [Saha et al., 2021]. The Grover coin provides a simple and highly symmetric choice for the internal dynamics of the walk. Different coin parametrisations would, in general, lead to distinct effective behaviours in the continuum limit. In the present work, we focus on the symmetric Grover-type coin in order to preserve translational invariance and maintain a balanced coupling between all internal states. This choice ensures that any emergent large scale dynamics arise solely from the interplay between coherent evolution and decoherence, rather than from asymmetries in the underlying coin operation. Nonetheless, the symmetry can be intentionally broken if one wishes to introduce directional bias or model anisotropic effects.
@@ 2.1 Fourier Basis
In order to analyse the continuum limit, it is convenient to work in the momentum representation, where the shift operator becomes diagonal and separable from the coin operator [Ambainis et al., 2001]. We therefore first transform both operators to Fourier space, before expressing them in the $\mathcal{SU}(3)$ Gell–Mann basis used for the continuum expansion. To obtain the momentum–space form of the walk, we apply the spatial Fourier transform:
@@
$\tilde{S}(k)=e^{-ik\Delta_{x}}|L\rangle\langle L|\;+\;|S\rangle\langle S|\;+\;e^{+ik\Delta_{x}}|R\rangle\langle R|,$ (8)
@@
$\tilde{\boldsymbol{\psi}}(k,t{+}1)=\mathbf{\tilde{S}}(k)\,\mathbf{\hat{C}}_{G}\,\tilde{\boldsymbol{\psi}}(k,t).$ (9)
@@
$\tilde{S}(k)=\begin{pmatrix}e^{-ik\Delta_{x}}&0&0\\[4.0pt] 0&1&0\\[4.0pt] 0&0&e^{ik\Delta_{x}},\end{pmatrix}$ (10)
which shows explicitly that the $|R\rangle$ component shifts forward, the $|L\rangle$ component shifts backward, and the $|S\rangle$ component remains localised.
The Grover coin, given by equation (3) admits a simple spectral decomposition, which enables us to parameterise it as the exponential of a generator. We define $G=\mathbb{I}_{3}-|u\rangle\langle u|$ with $|u\rangle$ being defined in equation (4), then
@@
A convenient parametrisation is given by
$\mathbf{\hat{C}}(\theta)=e^{-i\theta G}=e^{-i\theta(\mathbb{I}_{3}-|u\rangle\langle u|)}=|u\rangle\langle u|+e^{-i\theta}\,(\mathbb{I}_{3}-|u\rangle\langle u|).$ (12)
In this representation, the eigenstate corresponding to the uniform vector $|u\rangle$ has eigenvalue +1+1, while the two orthogonal Fourier modes acquire a common phase $e^{-i\theta}$. Thus, the coin operator $\hat{\mathbf{C}}(\theta)$ forms a smooth one parameter family that interpolates around the Grover operator. In particular, the standard Grover coin is recovered at the special value $\theta=\pi$, so that $\hat{\mathbf{C}}_{G}=\hat{\mathbf{C}}(\pi)$. This parametrisation introduces a tunable coin angle that controls the relative phase between the symmetric mode and the two transverse modes, and is therefore the natural small parameter for the continuum expansion. The parameter $\theta$ controls the strength of mixing between the three coin states. In the continuum limit this coupling plays the role of an effective mass term. Small $\theta$ yields nearly ballistic propagation, while larger values increase local mixing and open a mass-like gap in the dispersion. The full walk operator in momentum space can be written as:
$\mathbf{\tilde{U}}(k,\theta)=\mathbf{\tilde{S}}(k)\,\mathbf{\mathbf{\hat{C}}}(\theta)$ (13)
@@ 2.2 Gell Mann Representation
We now express $\tilde{S}(k)$ and $\mathbf{\hat{C}}(\theta)$ in the $\mathcal{SU}$(3) Gell Mann basis. This representation makes the underlying generators explicit and is the natural starting point for performing the continuum limit expansion of the walk operator in section 3.
For two state quantum walks, the two-dimensional coin operators can be decomposed in the Pauli basis, any $\mathcal{SU}$(2) operator can be written as a linear combination of Pauli matrices. In the lazy quantum walk, the coin space is three dimensional, so the natural generalisation is to use the eight Gell Mann matrices $\{\lambda_{i}\}_{i=1}^{8}$, traceless Hermitian operators, which, together with the identity $\mathbb{I}_{3}$, span the space of $3\times 3$ Hermitian operators [Wilson, 2024]. The explicit form of each Gell Mann matrix can be found in Appendix A. They satisfy the orthogonality relation $\mathrm{Tr}(\lambda_{i}\lambda_{j})=2\delta_{ij}$, so that any $\mathcal{SU}(3)$ operator $A\in\mathbb{C}^{3\times 3}$ can be expanded as
@@
Having expressed both the shift and coin in a separable form in Fourier space, we now rewrite these operators in the $\mathcal{SU}(3)$ Gell Mann basis. We first define the $\mathcal{SU}(3)$ generator:
$J_{z}^{(3)}\;=\;\tfrac{1}{2}\lambda_{3}+\tfrac{\sqrt{3}}{.}{2}\lambda_{8}\;=\;\mathrm{diag}(1,0,-1)$ (15)
@@
$\mathbf{\tilde{S}}(k)=\exp\!\bigl[-\,i\,\Delta_{x}k\,(\mathbb{I}_{x}\otimes J_{z}^{(3)})\bigr].$ (16)
@@
which mixes the three internal states.
@@ 2.3 Decoherence in Quantum Walks
Decoherence is usually described as a non-unitary evolution using the density matrix formalism to describe the mixed states that arise. Simple decoherence models for quantum walks were introduced in Kendon and Tregenna [2003]. Here, we use a similar simple Lindblad master equation to examine the basic features of decoherence in the lazy quantum walk, providing a baseline against which more complex decoherence mechanisms can be compared. This approach assumes a Markovian open system description, in which the environment has a short correlation time and no memory effects are retained. The Lindblad master equations provides dynamics that are completely positive and trace preserving, ensuring that the density operator remains physical at all times, regardless of the step size, as we take the limits [Annabestani et al., 2010]. The coupling to the environment is treated phenomenologically through a single decoherence strength $\gamma$, without specifying a microscopic bath model or spectral density. Initial system-environment correlations are neglected, and the system follows a weak decoherence regime [Breuer and Petruccione, 2002].
All the operators act in the same Hilbert space $\mathcal{H}_{xc}=\mathcal{H}_{x}\otimes\mathcal{H}_{c}$ as the unitary quantum walk. The system is described by a time dependent density operator, expressed in the computational basis as
$\boldsymbol{\hat{\rho}}(t)=\sum_{x,c}\sum_{x^{\prime},c^{\prime}}\rho_{x,c}^{x^{\prime},c^{\prime}}(t)|x,c\rangle\langle x^{\prime},c^{\prime}|.$ (18)
@@
$\boldsymbol{\hat{\rho}}(t+\Delta_{t})=(1-\gamma)\hat{\mathbf{U}}\boldsymbol{\hat{\rho}}(t)\hat{\mathbf{U}}^{\dagger}+\gamma\sum_{j}\mathbb{P}_{j}\hat{\mathbf{U}}\boldsymbol{\hat{\rho}}(t)\hat{\mathbf{U}}^{\dagger}\mathbb{P}_{j}^{\dagger},$ (19)
where $\gamma\in[0,1]$ is the probability of decoherence per full step, and the operator $\tilde{\mathbf{U}}$ is the full unitary evolution operator. This decoherence model was reviewed in Kendon [2007] and evolved numerically for various choices of $\mathbb{P}_{j}$. This is the infinite temperature Lindblad form, whose stationary state is the maximally mixed state and treats all energy levels symmetrically.
@@
The decoherence is modelled as projective noise using jump operators constructed as
@@
Figure 2 compares the effects of coin and spatial dephasing on the lazy DTQW for several values of the decoherence strength $\gamma$ and for different coin angles $\theta$. Coin dephasing, shown in panels (a,c), selectively damps the off-diagonal coherences between the internal coin states $\{|L\rangle,|S\rangle,|R\rangle\}$ while leaving spatial coherences intact. In contrast, spatial dephasing, shown in panels (b,d), suppresses off-diagonal density matrix elements $\rho(x,x^{\prime})$ for $x\neq x^{\prime}$. As $\gamma$ increases, the ballistic side peaks collapse more rapidly than in the coin dephased case and the probability distribution evolves more rapidly toward a broad, approximately Gaussian profile centred at the origin. This transition occurs already at relatively small values of $\gamma$, reflecting the sensitivity of ballistic transport to spatial coherence. For the spatially dephased case, the probability distribution seems to flatten out at low $\gamma$ values before tending towards the Gaussian profile. This is consistent with the behaviour of of the two state quantum walk under projective decoherence. In the limit $\gamma\to 1$, both decoherence mechanisms yield a similar Gaussian distribution, corresponding to a fully classical diffusive regime. Comparing different coin angles further reveals the role of the coin in controlling transport and localisation. For $\theta=\pi$, the unitary walk exhibits a pronounced central peak, indicating that the coin strongly favours population of the rest state and suppresses propagation. For $\theta=\pi/2$, the coin induces stronger mixing between the propagating and non-propagating components, resulting in a less pronounced central peak and a higher probability of reaching the ballistic fronts. This behaviour persists under both decoherence mechanisms, although spatial dephasing again accelerates the transition toward a Gaussian profile. Overall, increasing $\theta$ modifies the internal coin mixing and interference structure of the walk, leading to systematically reduced spatial spreading and enhanced localisation near the origin.